OpenStudy (anonymous):

Please help! If s(x) = x – 7 and t(x) = 4x2 – x + 3, which expression is equivalent to (t*s)(x) ?

OpenStudy (anonymous):

This is what i think the answer is :(t o s)(x) = t(s(x)) s(x) = x - 7 t(x - 7) = 4(x-7)^2 - (x - 7) + 3

OpenStudy (anonymous):

Hope that helps

OpenStudy (anonymous):

Thanks!